OpenStudy (anonymous):

Find the derivative of y with respect to x y= ln (1-x)/((x+2)^5)

OpenStudy (anonymous):

anyone please!!

OpenStudy (anonymous):

\[\left(\frac{a[x]}{b[x]}\right)'=\frac{a'[x]}{b[x]}-\frac{a[x] b'[x]}{b[x]^2} \]where a(x)=Log(1 - x) and b(x)=(x+2)^5

OpenStudy (anonymous):

Is just the answer satisfactory?

OpenStudy (anonymous):

hmm.

OpenStudy (anonymous):

i dont get what the answer is.

OpenStudy (anonymous):

\[-\frac{1}{(1-x) (x+2)^5}-\frac{5 \log (1-x)}{(x+2)^6} \]

OpenStudy (anonymous):

i got ln (1-x)(5(x+2)^4 all over (x+2)^10

OpenStudy (anonymous):

i dont have that as a choice.

OpenStudy (anonymous):

try separating expanding the logarthm first: \[\large y=ln\frac{1-x}{(x+2)^5}=ln(1-x)-5ln(x+2) \]

OpenStudy (anonymous):

what happened to the so we dont need to make it 5 ln (x+2)^4?

OpenStudy (anonymous):

i think finding the derivative will be easier from here if you use y'/y

OpenStudy (anonymous):

@MegMegs4 , no... it's the property of logs...\(\large log_bM^n=nlog_bM \)

OpenStudy (anonymous):

oh ok i think thats where i was making my mistake.

OpenStudy (anonymous):

so i got.... -1/(1-x) - 5/(x+2)

OpenStudy (anonymous):

that looks good...:)

OpenStudy (anonymous):

ok but it isnt one of my answer choices.

OpenStudy (anonymous):

subtract them....

OpenStudy (anonymous):

my choices are a. (x+2)^5 / (1-x) b. ln (6x-7)/(x+2)^6 c. 4x-7 / (x+2)(1-x) d. 4x-7 /(x+2)^6

OpenStudy (anonymous):

im trying it now.

OpenStudy (anonymous):

yea... it looks as though they did subtract them..

OpenStudy (anonymous):

so the answer is c!! thank you . you are so smart!!

OpenStudy (anonymous):

that's what i got..

OpenStudy (anonymous):

thanks... yw..:)

OpenStudy (anonymous):

@dpaInc Isn't the problem equation the following ?\[y=\frac{\log (1-x)}{(x+2)^5} \]